# biblio.bib

### v3

@@ 4 Extension to labelled DAGs
In this section we demonstrate how our decomposition scheme can be applied to the classical model of labelled DAGs to obtain new recurrences on known sequences. In the 1990s, Gessel [Gessel1996] already obtained equations allowing to count labelled DAGs by vertices, edges, and sources (and also sinks actually) using a generating functions approach. These equations involve the inclusion-exclusion principle which has one drawback: they are usually not amenable to efficient random generation. The reason for this is that subtractions translate into rejections in the recursive algorithm, that are here too costly to be usable. In the present paper, we derive new recurrences with a combinatorial meaning and that do not involve the inclusion-exclusion principle. As a consequence, we can obtain an efficient random sampler of DAGs with full control over the number of vertices, sources, and edges.
As in Section 2, we present here a slight generalisation of the formula given in [GPV2021, §4] allowing to capture various classes of labelled DAGs. We omit the proofs here as they follow a straightforward adaptation of the arguments given in [GPV2021] and the recursive method.
@@ 4.1 Recursive decomposition
Let $\mathcal{P}$ be any subset of $\mathbb{N}$ and let $A_{n,m,k}^{\mathcal{P}}$ denote the number of labelled DAGs with $m$ edges and $n$ vertices including $k$ sources, and in which every vertex except the (first) sink has out-degree in $\mathcal{P}$. The number of such DAGs with a distinguished (or pointed) source is given by $k\cdot A_{n,m,k}^{\mathcal{P}}$ since any of the $k$ sources may be distinguished. Let $D$ denote one such DAG and let $v$ denote its distinguished source. Removing the distinguished source in $D$ and decrementing the labels of the vertices of higher label than $v$ by one yields a regular vertex-labelled DAG $D^{\prime}$ with $n-1$ vertices. Moreover, the three pieces of information that are necessary to reconstruct the source are the following:
@@
$kA_{n,m,k}^{\mathcal{P}}$ =$=\begin{cases}n\sum_{p\in\mathcal{P}\cap\llbracket 0;n-k\rrbracket}\sum_{i=0}^{p}A_{n-1,m-p,k-1+p-i}^{\mathcal{P}}\binom{n-k-p+i}{i}\binom{k-1+p-i}{p-i}&\text{if~{}$1\leq k$}\\ 0&\text{otherwise.}\end{cases}$
@@
When $\mathcal{P}=\mathbb{N}$, we recover the sequence counting all labelled DAGs, known as A003024 in the OEIS and first enumerated in [Robinson1970, Stanley1973, Robinson1973]. For $\mathcal{P}=\mathbb{N}^{\star}$ and with $k=1$, we find the number of labelled DAGs with a single source and a single sink, known to Gessel in [Gessel1995, Gessel1996] and stored at A165950 in the OEIS.
@@ 4.2 Random generation
         $D^{\prime}$leftarrow unifDAGP($n-1,m-p,k-1+p-i$)
         $I$leftarrow a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
         $S$leftarrow a uniform subset of size $(p-i)$ the sources of $D^{\prime}$
         $\ell$leftarrow Unif($\llbracket 1;n\rrbracket$)
@@ 5.1 The encoding
The decomposition scheme described in Section 2 corresponds to a traversal of the DOAG. This traversal induces a labelling of the vertices from $1$ to $n$, which allows us to associate the vertices of the graph to these integers in a canonical way. We then consider its transition matrix using these labels as indices. Usually, the transition matrix of a directed graph $D$ is defined as the matrix ${(a_{i,j})}_{1\leq i,j\leq n}$ such that $a_{i,j}$ is $1$ if there is an edge from vertex $i$ to vertex $j$ in $D$, and 00 otherwise. This representation encodes the set of the edges of a DAG but not the edge ordering of DOAGs. In order to take this ordering into account, we use a slightly different encoding.
@@
[…]
@@
  • for all $11\leq i\leq n-$ 1, the sub-row ${(a_{i,j})}_{i<j\leq n}$ is a variation (of size $n-i$).
From an enumerative point of view, a variation matrix can be seen as a sequence of variations $(v_{1},v_{2},\ldots,v_{n-1})$ where for all $11\leq i\leq n-$ 1, the variation $v_{i}$ has size $i$.
@@
Another important remark is that, when several underlined cells occur on the same row $i$ in the matrix, they correspond to several sources that are discovered at the same decomposition step of the DOAG (upon removing the same source). Recall that the decomposition
@@
algorithm sorts the labels of these new sources by following the total order of the outgoing edges of vertex $i$. This implies that the underlined entries within the same row have to be increasingly sorted (from left to right). For instance, observe that there are three consecutive underlined cells in the first row of the matrix in Figure 6. Indeed, when removing the first source of the DOAG on the left, we uncover three new sources which are respectively in first, second and fourth position in the outgoing edges order of the removed source.
@@
All labelled transition matrices of DOAGs are variation matrices. Furthermore, let $A={(a_{i,j})}_{1\leq i,j\leq n}$ be a variation matrix, and for all $j\in\llbracket 1;n\rrbracket$, let $b_{j}$ denote the largest $i\leq n$ such that $a_{i,j}>0$ if such an index exists and 00 otherwise. Then, $A$ is the labelled transition matrix of some DOAG if and only if the two following properties hold:
@@ 6 Asymptotic results
[…]
@@
[…]
@@ 6.1 First bounds on the number of DOAGs with $\boldsymbol{n}$ vertices
[…]
Lemma 2 (Asymptotic estimation of the super-factorial [Barnes1900]).
@@
[…]
@@
Let $0\leq p\leq n$, a variation of size $n$ containing exactly $p$ zeros is the interleaving of a permutation of size $(n-p)$ with an array of zeros of size $p$. As a consequence
@@
In a labelled transition matrix with positive values on the super-diagonal, the $i$-th row can be seen as a variation of size $(n-i)$ that does not start with a zero. Moreover, the number of variations of size $n$ starting with a zero is actually the number of variations of size $(n-1)$ so that the number of possibilities for the $i$-th row of the matrix we count here is $(v_{n-i}-v_{n-i-1})$. In addition, by Lemma 3, we also have that
@@
Note that when $i=n-1$, we have $v_{n-i}-v_{n-i-1}=v_{1}-v_{0}=1$. Indeed, the row of index $i=n-1$ contains only the number $1$ in the super diagonal since at the last step of the decomposition, we have two connected vertices and there is only one such DOAG. Setting aside this special case, which does not contribute to the product, we get the following lower bound for $D_{n}$:
[…]
@@
[…]
@@
In order to obtain this lower bound, we count the number of valid labelled transition matrices such that all but exactly one of the cells on the super-diagonal have non-zero values. Furthermore, in order to avoid having to deal with border cases, we assume that the unique zero value on the super-diagonal appears between $i=2$ and $i=n-2$. Let $22\leq i\leq n-$ 2 and let us consider those matrices ${(a_{p,q})}_{1\leq p,q\leq n}$ such that $a_{i,i+1}=0$. Those matrices are illustrated in Figure 10
@@
    On row $i-1$, the two first cells on the right of the diagonal ($a_{i-1,i}$ and $a_{i-1,i+1}$) must have positive values and must be in increasing order. In the case of $a_{i-1,i}$, this is because it is on the super diagonal. And for $a_{i-1,i+1}$, this is because $a_{i,i+1}=0$: since it is above this cell, and since the cell $a_{i-1,1}$ on its left is non-zero, it must be non-zero. Otherwise the condition from Theorem 3 are violated. In terms of DOAGs, this means that the vertex $i-1$ produces two new sources when it is removed but vertex $i$ produces none.
@@
    On row $i$, any variations of size $(n-i)$ starting by a zero is allowed.
@@
[…]
@@ 6.2 Obtaining the polynomial term by bootstrapping
In this case, the first source is not connected to the second vertex and the matrix has thus more than one source. The first row of such a matrix is a variation of size $(n-2)$ and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size $(n-1)$, the DOAG obtained by removing the first source. However, it is important to note that not all combinations of a size-$(n-2)$ variation and a size-$(n-1)$ matrix yield a valid size-$n$ labelled transition matrix. For instance, a variation of the form $v=(0,1,0,2,\ldots)$ and a lower matrix with at least three sources cannot be obtained together as they would violate the constraints of Theorem 3.
@@
In this case, the first row is a variation of size $(n-1)$ starting by a positive value, and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size $(n-1)$ with exactly one source, again obtained by removing the first source. This second fact is a direct consequence of $a_{2,3}>0$. Here, all such pairs can be obtained.
@@
In this case the lower part ${(a_{i,j})}_{3\leq i,j\leq n}$ encodes a DOAG of size $n-2$ with exactly one source, the first row is necessarily a variation of size $(n-1)$, starting with two positive increasing values, and the second row is a variation of size $(n-2)$ starting by a zero. Here again this decomposition is exact: all such triplets can be obtained here.
@@
Finally, this case captures all the remaining matrices. The first row is a variation of size $(n-1)$, the second and third rows are variations of sizes $(n-2)$ and $(n-3)$ starting with a zero, and the lower part ${(a_{i,j})}_{4\leq i,j\leq n}$ encodes a size-$(n-3)$ DOAG. Of course, not all such quadruples can be obtained, but this over-approximation will be enough for our proof.
@@
Figure 11: Decomposition of DOAG labelled transition matrices based or their content near the top of the diagonal. The symbols $\mathcal{D}$ and ${\mathcal{D}}^{\star}$ respectively represent the set of all possible DOAG labelled transition matrices the set of all of those matrices such that $a_{1,2}>0$. The stars ()($\star$)( ⋆ ) represent strictly positive values.
@@
Note that the big-$O$ term comes from the fact that not all pairs made of a size-$(n-2)$ variation and a size-$(n-1)$ labelled transition matrix can be obtained this way, as discussed in the first case above. We could actually have written $0\leq D_{n}-D_{n}^{\star}\leq v_{n-2}D_{n-1}$.
@@
[…]
@@
[…]

### v4

@@ 4 Extension to labelled DAGs
In this section we demonstrate how our decomposition scheme can be applied to the classical model of labelled DAGs to obtain new recurrences on known sequences. In the 1990s, Gessel [19] already obtained equations allowing to count labelled DAGs by vertices, edges, and sources (and also sinks actually) using a generating functions approach. These equations involve the inclusion-exclusion principle which has one drawback: they are usually not amenable to efficient random generation. The reason for this is that subtractions translate into rejections in the recursive algorithm, that are here too costly to be usable. In the present paper, we derive new recurrences with a combinatorial meaning and that do not involve the inclusion-exclusion principle. As a consequence, we can obtain an efficient random sampler of DAGs with full control over the number of vertices, sources, and edges.
As in Section 2, we present here a slight generalisation of the formula given in [18, §4] allowing to capture various classes of labelled DAGs. We omit the proofs here as they follow a straightforward adaptation of the arguments given in [18] and the recursive method.
@@ 4.1 Recursive decomposition
Let $\mathcal{P}$ be any subset of $\mathbb{N}$ and let $A_{n,m,k}^{\mathcal{P}}$ denote the number of labelled DAGs with $m$ edges and $n$ vertices including $k$ sources, and in which every vertex except the (first) sink has out-degree in $\mathcal{P}$. The number of such DAGs with a distinguished (or pointed) source is given by $k\cdot A_{n,m,k}^{\mathcal{P}}$ since any of the $k$ sources may be distinguished. Let $D$ denote one such DAG and let $v$ denote its distinguished source. Removing the distinguished source in $D$ and decrementing the labels of the vertices of higher label than $v$ by one yields a regular vertex-labelled DAG $D^{\prime}$ with n1n-1 vertices. Moreover, the three pieces of information that are necessary to reconstruct the source are the following:
@@
[…]
@@
When $\mathcal{P}=\mathbb{N}$, we recover the sequence counting all labelled DAGs, known as A003024 in the OEIS and first enumerated in [38, 41, 36]. For $\mathcal{P}=\mathbb{N}^{\star}$ and with $k=1$, we find the number of labelled DAGs with a single source and a single sink, known to Gessel in [20, 19] and stored at A165950 in the OEIS.
@@ 4.2 Random generation
   $D^{\prime}$ $\leftarrow$ unifDAGP(n1,mp,k1+pin-1,m-p,k-1+p-i)
   II $\leftarrow$ a uniform subset of size $i$ of the inner vertices of $D^{\prime}$
   $S$ $\leftarrow$ a uniform subset of size (pi)(p-i) the sources of $D^{\prime}$
   $\ell$ $\leftarrow$ Unif($\llbracket 1;n\rrbracket$)
@@ 5.1 The encoding
The decomposition scheme described in Section 2 corresponds to a traversal of the DOAG. This traversal induces a labelling of the vertices from $1$ to $n$, which allows us to associate the vertices of the graph to these integers in a canonical way. We then consider its transition matrix using these labels as indices. Usually, the transition matrix of a directed graph $D$ is defined as the matrix ${(a_{i,j})}_{1\leq i,j\leq n}$ such that $a_{i,j}$ is $1$ if there is an edge from vertex $i$ to vertex $j$ in $D$, and 0 otherwise. This representation encodes the set of the edges of a DAG but not the edge ordering of DOAGs. In order to take this ordering into account, we use a slightly different encoding.
@@
For instance, the sequence $(6,2,3,0,0,1,4,0,5)$ is a variation of size $9$ but the sequences $(1,0,3)$ and $(1,0,3,2,3)$ are not variations because the number $2$ is missing in the first one and the second contains two occurrences of the number $3$. Variations can also be defined as interleavings of a permutation with a sequence of zeros. One of the earliest references to these objects dates back to 1659 in Izquierdo’s Pharus Scientiarum [25, Disputatio 29]. They also appear in Stanley’s book as the second entry of his Twelvefold Way [42, page 79], a collection of twelve basic but fundamental counting problems. Knuth gives a few ancient references on this topic in [29] and in an quote (without reference) that can be found on the OEIS page of variations at A007526. Variations are relevant to our problem as they naturally appear as rows of the labelled transition matrices defined in this section. Some of their combinatorial properties will be exhibited in the next section.
@@
  • for all $1\leq i\leq n-1$, the sub-row ${(a_{i,j})}_{i<j\leq n}$ is a variation (of size nin-i).
From an enumerative point of view, a variation matrix can be seen as a sequence of variations $(v_{1},v_{2},\ldots,v_{n-1})$ where for all $1\leq i\leq n-1$, the variation $v_{i}$ has size $i$.
@@
Another important remark is that, when several underlined cells occur on the same row $i$ in the matrix, they correspond to several sources that are discovered at the same decomposition step of the DOAG (upon removing the same source).
@@
Recall that the decomposition algorithm sorts the labels of these new sources by following the total order of the outgoing edges of vertex $i$. This implies that the underlined entries within the same row have to be increasingly sorted (from left to right). For instance, observe that there are three consecutive underlined cells in the first row of the matrix in Figure 6. Indeed, when removing the first source of the DOAG on the left, we uncover three new sources which are respectively in first, second and fourth position in the outgoing edges order of the removed source.
@@
All labelled transition matrices of DOAGs are variation matrices. Furthermore, let $A={(a_{i,j})}_{1\leq i,j\leq n}$ be a variation matrix, and for all $j\in\llbracket 1;n\rrbracket$, let $b_{j}$ denote the largest $i\leq n$ such that $a_{i,j}>0$ if such an index exists and 0 otherwise. Then, $A$ is the labelled transition matrix of some DOAG if and only if the two following properties hold:
@@ 6 Asymptotic results
$D_{n}$ $=\left(1+O\!\left(n^{-1}\right)\right)D_{n}^{\star}=\frac{c}{\sqrt{n}}\ e^{n-1}\ \text{¡}{(n-1)}!\left(1+O\!\left(n^{-1}\right)\right)$
@@
where $\text{¡}{k}!=\prod_{i=0}^{k}{i!}$ denotes the super factorial of $k$.
@@ 6.1 First bounds on the number of DOAGs with $\boldsymbol{n}$ vertices
$D_{n}\leq\text{¡}{(n-1)}!e^{n-1}$
where $\text{¡}{k}!=\prod_{i=0}^{k}{i!}$ denotes the super factorial of $k$.
The term “super factorial” seems to have been coined by Sloane and Plouffe in [39, page 228] but this sequence had been studied before that, in 1900, by Barnes [2] as the integer values of the “G-function”. In fact, if $G(z)$ denotes the complex-valued G-function of Barnes, we have the identity $G(n+2)=\text{¡}{n}!$ for all integer $n$. Barnes also gives the following equivalent.
Lemma 2 (Asymptotic estimation of the super-factorial [2]).
@@
$\text{¡}{(n-1)}!=G(n+1)\sim e^{\zeta^{\prime}(-1)}\cdot n^{-1/12}\cdot{\big(\sqrt{2\pi}\big)}^{n}\cdot e^{-\frac{3}{4}n^{2}}\cdot n^{n^{2}/2}$
@@
Let $0\leq p\leq n$, a variation of size $n$ containing exactly $p$ zeros is the interleaving of a permutation of size (np)(n-p) with an array of zeros of size $p$. As a consequence
@@
In a labelled transition matrix with positive values on the super-diagonal, the $i$-th row can be seen as a variation of size (ni)(n-i) that does not start with a zero. Moreover, the number of variations of size $n$ starting with a zero is actually the number of variations of size (n1)(n-1) so that the number of possibilities for the $i$-th row of the matrix we count here is $(v_{n-i}-v_{n-i-1})$. In addition, by Lemma 3, we also have that
@@
Note that when i=n1i=n-1, we have $v_{n-i}-v_{n-i-1}=v_{1}-v_{0}=1$. Indeed, the row of index i=n1i=n-1 contains only the number $1$ in the super diagonal since at the last step of the decomposition, we have two connected vertices and there is only one such DOAG. Setting aside this special case, which does not contribute to the product, we get the following lower bound for $D_{n}$:
$D_{n}\geq\prod_{i=1}^{n-2}(v_{n-i}-v_{n-i-1})=e^{n-2}\text{¡}{(n-1)}!\prod_{i=1}^{n-2}\frac{n-1-i}{n-i}\prod_{i=1}^{n-1}\left(1+O\left(\frac{1}{(n-i)!}\right)\right)$ (5)
@@
Although they are not precise enough to obtain an asymptotic equivalent for the sequence $D_{n}$, these two bounds already give us a good understanding of the behaviour of $D_{n}$. First of all, they let appear a “dominant” term of the form $\text{¡}{(n-1)}!e^{n-1}$, which is uncommon in combinatorial enumeration. And second, it tells us we only make a relative error of the order of $O(n)$ when approximating $D_{n}$ by $\text{¡}{(n-1)}!\cdot e^{n-1}$. We will prove an asymptotic equivalent for the remaining polynomial term, but in order to obtain this, we first need to slightly refine our lower bound so that the “interval” between our two bounds is a little narrower than $O\!\left(n\right)$.
@@
In order to obtain this lower bound, we count the number of valid labelled transition matrices such that all but exactly one of the cells on the super-diagonal have non-zero values. Furthermore, in order to avoid having to deal with border cases, we assume that the unique zero value on the super-diagonal appears between $i=2$ and i=n2i=n-2. Let $2\leq i\leq n-2$ and let us consider those matrices ${(a_{p,q})}_{1\leq p,q\leq n}$ such that $a_{i,i+1}=0$. Those matrices are illustrated in Figure 10
@@
    On row i1i-1, the two first cells on the right of the diagonal ($a_{i-1,i}$ and $a_{i-1,i+1}$) must have positive values and must be in increasing order. In the case of $a_{i-1,i}$, this is because it is on the super diagonal. And for $a_{i-1,i+1}$, this is because $a_{i,i+1}=0$: since it is above this cell, and since the cell $a_{i-1,1}$ on its left is non-zero, it must be non-zero. Otherwise the condition from Theorem 3 are violated. In terms of DOAGs, this means that the vertex i1i-1 produces two new sources when it is removed but vertex $i$ produces none.
@@
    On row $i$, any variations of size (ni)(n-i) starting by a zero is allowed.
@@
In addition, we know from the proof of Lemma 4 that the product in equation (6) is equivalent to $\frac{c}{n}e^{n-1}\text{¡}{(n-1)}!$ for some constant $c$, which allows to conclude. ∎
@@ 6.2 Obtaining the polynomial term by bootstrapping
In this case, the first source is not connected to the second vertex and the matrix has thus more than one source. The first row of such a matrix is a variation of size (n2)(n-2) and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size (n1)(n-1), the DOAG obtained by removing the first source. However, it is important to note that not all combinations of a size-(n2)(n-2) variation and a size-(n1)(n-1) matrix yield a valid size-$n$ labelled transition matrix. For instance, a variation of the form $v=(0,1,0,2,\ldots)$ and a lower matrix with at least three sources cannot be obtained together as they would violate the constraints of Theorem 3.
@@
In this case, the first row is a variation of size (n1)(n-1) starting by a positive value, and the lower part ${(a_{i,j})}_{2\leq i,j\leq n}$ encodes a DOAG of size (n1)(n-1) with exactly one source, again obtained by removing the first source. This second fact is a direct consequence of $a_{2,3}>0$. Here, all such pairs can be obtained.
@@
In this case the lower part ${(a_{i,j})}_{3\leq i,j\leq n}$ encodes a DOAG of size n2n-2 with exactly one source, the first row is necessarily a variation of size (n1)(n-1), starting with two positive increasing values, and the second row is a variation of size (n2)(n-2) starting by a zero. Here again this decomposition is exact: all such triplets can be obtained here.
@@
Finally, this case captures all the remaining matrices. The first row is a variation of size (n1)(n-1), the second and third rows are variations of sizes (n2)(n-2) and (n3)(n-3) starting with a zero, and the lower part ${(a_{i,j})}_{4\leq i,j\leq n}$ encodes a size-(n3)(n-3) DOAG. Of course, not all such quadruples can be obtained, but this over-approximation will be enough for our proof.
@@
Figure 11: Decomposition of DOAG labelled transition matrices based or their content near the top of the diagonal. The symbols $\mathcal{D}$ and $\mathcal{D}$$^{\star}$ respectively represent the set of all possible DOAG labelled transition matrices the set of all of those matrices such that $a_{1,2}>0$. The stars $(\star)$ represent strictly positive values.
@@
Note that the big-$O$ term comes from the fact that not all pairs made of a size-(n2)(n-2) variation and a size-(n1)(n-1) labelled transition matrix can be obtained this way, as discussed in the first case above. We could actually have written $0\leq D_{n}-D_{n}^{\star}\leq v_{n-2}D_{n-1}$.
@@
Let us introduce the polynomial term $P_{n}^{\star}$ of $D_{n}^{\star}$ defined by $P_{n}^{\star}=D_{n}^{\star}/e^{n-1}\text{¡}{(n-1)}!$. By normalising equation (8) and using equation (7) we have
@@
In other words, we have that $P_{n}$ and $P_{n}^{\star}$ are equivalent. Then, by normalising equation (9) by $e^{n-1}\text{¡}{(n-1)}!$, we obtain the following asymptotic expansion